Bennett purchased some beef and some chicken for a family barbecue. The beef cost $5.90 per pound and the chicken cost $4.30 per pound. He brought a total of 19 lb of meat and spent $96.10. How much of each type of meat did he purchase?
He brought _____ lb of beef.

Answer by Boreal(15235)

x=beef
c=chicken
5.90x+4.30(19-x)=96.10, because x+c=19
5.90x+81.70-4.30x=96.10
1.6x=14.40
x=9 pounds beef $53.10
19=x=10 pounds chicken $43.00